### v2

# On the Use of Hierarchical Vision Foundation Models for Low-Cost Human Mesh Recovery and Pose Estimation

Abstract
In this work, we aim to develop simple and efficient models for human mesh recovery (HMR) and its predecessor task, human pose estimation (HPE). State-of-the-art HMR methods, such as HMR2.0 and its successors, rely on large, non-hierarchical vision transformers as encoders, which are inherited from the corresponding HPE models like ViTPose. To establish baselines across varying computational budgets, we first construct three lightweight HMR2.0 variants by adapting the corresponding ViTPose models. In addition, we propose leveraging the early stages of hierarchical vision foundation models (VFMs), including Swin Transformer, GroupMixFormer, and VMamba, as encoders. This design is motivated by the observation that intermediate stages of hierarchical VFMs produce feature maps with resolutions comparable to or higher than those of non-hierarchical counterparts. We conduct a comprehensive evaluation of 27 hierarchical-VFM-based HMR and HPE models, demonstrating that using only the first two or three stages achieves performance on par with full-stage models. Moreover, we show that the resulting truncated models exhibit better trade-offs between accuracy and computational efficiency compared to existing lightweight alternatives. The source code is available at https://github.com/nttcom/TruncHierVFM.
[Uncaptioned image][Uncaptioned image]
Figure 1: (Left): 3D pose estimation results (PA-MPJPE; lower is better) of human mesh recovery (HMR) models on the Human3.6M dataset [ionescu+2014tpami]. (Right): 2D pose estimation results (AP; higher is better) of human pose estimation (HPE) models on the COCO-val dataset [lin+2014eccv]. Circle size reflects model size. Gray circles indicate existing methods, while colored circles represent our proposed models, which leverage the early stages of hierarchical vision foundation models (VFMs) as encoders. These results demonstrate that our models tend to offer a more favorable trade-off between performance and computational cost (i.e., GFLOPs) compared to existing approaches.
1 Introduction
Human mesh recovery (HMR) plays a central role in a wide range of applications, including animation, virtual try-on, sports analytics, and human-computer interaction [zheng+2023acm, tian+2023tpami, gao+2025air, liu+2024neuro]. Over the past decade, this research field has witnessed remarkable progress, driven in part by vision foundation models (VFMs) [he+2016cvpr, wang+2021tpami, sun+2019cvpr, dosovitskiy+2020iclr]. While early HMR approaches [kanazawa+2018cvpr, zhang+2021iccv, kolotouros+2021iccv, li+2022eccv2, kocabas+2021iccv] primarily relied on convolutional neural networks (CNNs), recent state-of-the-art (SoTA) methods [cho+2022eccv, dou+2023iccv, lin+2021cvpr, lin+2021iccv, agarwal+2024eccvw, pang+2022neurips, cai+2024tpami, romain+2025cvprw, lin+2025cvpr, feng+2024cvpr] have increasingly adopted Transformer-based architectures [vaswani+2017nips]. Among them, HMR2.0 [goel+2023iccv] has garnered significant attention for its simplicity and strong performance. HMR2.0 and its successors [patel+20243dv, fiche+2025cvpr, saleem+2025aaai, prospero+2025cvprw, stathopoulos+2024cvpr, dwivedi+2024cvpr] employ a large non-hierarchical vision transformer (i.e., ViT-H [dosovitskiy+2020iclr]) as their encoder, which is inherited from the corresponding human pose estimation (HPE) model (i.e., ViTPose-H [xu+2022neurips]).
In general, even for HMR and HPE, large VFMs demand substantial computational resources, which can hinder their deployment in real-time or resource-constrained settings such as mobile devices or edge computing environments. With HMR2.0, a straightforward approach to alleviate this issue is to use smaller ViT variants (e.g., ViT-L, ViT-B, ViT-S) as encoders. Since this direction has not been explored in the literature, in this work we instantiate smaller variants of HMR2.0 as baselines (see §3.2.1 for details). Beyond this, to better balance performance and efficiency while preserving the architectural simplicity of HMR2.0, we investigate the use of hierarchical VFMs [liu+2021iccv, ge+2023arxiv, liu+2024neurips] as encoders for HMR and its predecessor, HPE. The key insight motivating our approach lies in the resolution characteristics of intermediate representations in hierarchical VFMs. They typically follow a four-stage structure, where the spatial resolution of feature maps are higher or the same with the consistent resolution seen in non-hierarchical VFMs. Therefore, if the intermediate outputs of pretrained hierarchical VFMs retain sufficient semantic richness and spatial detail, the latter stages of the original backbone can be removed. This allows for reductions in model size and computational cost without compromising architectural simplicity.
We conduct extensive experiments to validate this observation by instantiating HMR and HPE models within the HMR2.0 and ViTPose frameworks, using different stages of three hierarchical VFMs as encoders: Swin Transformer [liu+2021iccv], GroupMixFormer [ge+2024cvpr], and VMamba [liu+2024neurips]. In total, we instantiate 27 hierarchical-VFM-based HMR and HPE models. Our results consistently show that models using only the first two or three VFM stages as encoders achieve performance comparable to, and occasionally better than, their full four-stage counterparts. Moreover, these truncated models demonstrate a more favorable trade-off between accuracy and efficiency compared to existing lightweight approaches, including the small ViT-based variants of HMR2.0 and ViTPose. Our contributions are summarized as follows:
  • We instantiate lightweight variants of HMR2.0 via inheriting the encoders of ViTPose-{L,B,S}.
  • We investigate the use of hierarchical VFMs [liu+2021iccv, ge+2023arxiv, liu+2024neurips] as encoders within the HMR2.0 and ViTPose frameworks. Experimental results show that models utilizing only the first few VFM stages as their encoders achieve performance comparable to, and sometimes better than, their full four-stage counterparts for both HMR and HPE.
  • We further demonstrate that hierarchical-VFM-based HMR and HPE models achieve superior trade-offs between performance and efficiency compared to existing lightweight approaches, including those based on ViT.
(a) ViTPose [xu+2022neurips]
(b) HMR2.0 [goel+2023iccv]
Figure 2: Architectures of the baseline models for human pose estimation (HPE) and human mesh recovery (HMR).
2 Related Work
2.1 Human Mesh Recovery (HMR)
HMR has been extensively studied under a variety of problem settings, including multi-person HMR [romain+2025cvprw, liao+2024tvcg, sun+2021iccv, sun+2022cvpr, su+2025cvpr, fabien+2024eccv, sun+2024cvpr, choi+2022cvpr, zanfir+2018cvpr, zanfir+2018neurips, qiu+2023cvpr], multi-view HMR [bogo+eccv2016, zhang+2023iccv, zhu+2025aaai, li+2021wacv, pavlakos+2022cvpr, yu+2022neurips, jia+2023aaai, xie+2024icmr], video-based HMR [zhang+2025cvpr, choi+2021cvpr, kocabas+2020cvpr, wei+2022cvpr, kang+2025cvprw, shen+2024sa, yang+2023iclr], full-body HMR [pavlakos+2019cvpr, wang+2022cvpr, shen+2024acmmm, zhang+2023tpami, li+2025tpami, lin+2023cvpr, cai+2023neurips], privacy-preserving HMR [gao+2025air, wu+2025cvpr, ge+2024cvpr], and prompt-based HMR [wang+2025cvpr, lin+2025cvpr, feng+2024cvpr]. In this work, we focus on the most fundamental setting, image-based HMR, where the objective is to predict SMPL [loper+2015tog] parameters from a single cropped image of a person.
While early works employed CNNs such as ResNet [he+2016cvpr], HRNet [sun+2019cvpr, wang+2021tpami], and EfficientNet [tan+2019icml] as backbone encoders [fiche+2024eccv, le+2024cvpr, ma+2023cvpr, kanazawa+2018cvpr, zhang+2021iccv, kolotouros+2021iccv, shen+2025icml, sarandi+2024neurips, xiao+2024eccv, li+2022eccv2, kocabas+2021iccv, li+2024tip, xu+2024cvpr], they have been superseded by Transformer-based methods in recent years. For instance, METRO [lin+2021cvpr, cho+2022eccv], Mesh Graphormer [lin+2021iccv] and PointHMR [kim+2023cvpr] introduce hybrid CNN-Transformer architectures that leverage pretrained CNN features while capturing global dependencies via self-attention. On the other hand, several recent works [goel+2023iccv, lin+2023cvpr, cai+2024tpami, cai+2023neurips] utilize fully Transformer-based encoders. HMR2.0 [goel+2023iccv], along with its successors [patel+20243dv, fiche+2025cvpr, saleem+2025aaai, prospero+2025cvprw, stathopoulos+2024cvpr, dwivedi+2024cvpr], employs ViT-H as the encoder, initialized through pretraining on HPE tasks. SMPLer-X [cai+2023neurips], uses four ViTs to construct full-body HMR models with varying model sizes. More recently, prompt-based HMR approaches including ChatPose [feng+2024cvpr] and ChatHuman [lin+2025cvpr] leverage the ViT encoder from CLIP [radford+2021icml] to align visual and textual features.
Previous works have primarily explored the use of non-hierarchical transformers, ViT or DeiT, for the HMR task. In contrast, this study investigates the application of hierarchical VFMs, encompassing not only transformer-based architectures but also recently proposed state space models (SSMs). A notable exception in the literature is [pang+2022neurips], which integrates hierarchical transformers such as the Swin Transformer [liu+2021iccv, liu+2022cvpr] and Twins [chu+2021neurips] into the HMR framework of [kanazawa+2018cvpr]. DeFormer [yoshiyasu+2023cvpr] also explores the use of hierarchical transformers, such as the Swin Transformer with a Feature Pyramid Network and the Mix Transformer [cao+2022ijcai, xie+2021neurips], as backbones. Nevertheless, as discussed in §1, our work goes beyond simply applying hierarchical VFMs to HMR: We further investigate the use of only the initial stages of these hierarchical models as encoders, aiming to develop more efficient HMR architectures.
Several studies have aimed to develop efficient HMR models [zheng+2023cvpr, dou+2023iccv, agarwal+2024eccvw, wang+2025visapp]. For instance, CoarseMETRO [agarwal+2024eccvw] employs a coarse-to-fine strategy to reduce the computational burden of early transformer layers, while TORE [dou+2023iccv] accelerates HMR by pruning background tokens. POTTER [zheng+2023cvpr] integrates a high-resolution stream with a basic stream to recover more accurate human meshes while reducing memory usage and computational cost. Although effective, these methods often introduce additional architectural complexity. In contrast, our approach, leveraging the early stages of hierarchical VFMs, offers a simpler alternative that is potentially complementary to these techniques.
2.2 Human Pose Estimation (HPE)
Since the encoders in HMR methods are often inherited from those used in HPE models, another promising direction for developing efficient HMR is to adapt efficient HPE approaches. Several studies have explored this in the context of HPE. For example, MEMe [li+2022sensors], Lite-HRNet [yu+2021cvpr], and LitePose [wang+2022cvpr] focus on optimizing popular CNN backbones such as EfficientNet [tan+2019icml] and HRNet [sun+2019cvpr, wang+2021tpami] to improve the trade-off between accuracy and efficiency. DANet [luo+2021icme] proposes an improved multi-scale feature fusion strategy that eliminates the need for computationally expensive cascaded pyramid architectures. SimCC [li+2022eccv] reformulates HPE as two independent classification tasks for horizontal and vertical coordinates. Additionally, CNF [yang+2022pr] and ViPNAS [xu+2021cvpr] apply neural architecture search to automatically optimize network structures for improved efficiency.
All of the aforementioned approaches are based on CNN architectures. While CNNs can be considered a type of hierarchical VFM, in this work we focus on adapting more recent architectures, such as transformers and SSMs.
2.3 Vision Foundation Model (VFM)
VFMs can be broadly categorized into non-hierarchical models [dosovitskiy+2020iclr, touvron+2021icml, steiner+2022tmlr, beyer+2023cvpr, fang+2024ivc, tian+2023cvpr, wei+2023iccv, yang+2023iclr2, alkin+2025iclr, yang+2022pr] and hierarchical models [liu+2021iccv, fan+2021iccv, wang+2021iccv, zhang+2021iccv, li+2022cvpr, wang+2022cvmj, sun+2023tpami, zhang+2023iclr, ge+2023arxiv, liu+2024neurips]. Following the discussion above, this work focuses on recent hierarchical VFMs. Among them, we select Swin Transformer (Swin) [liu+2021iccv], GroupMixFormer (GMF), and VMamba (VM) [liu+2022cvpr] based on their strong performance, widespread recognition, and the availability of public code with pretrained weights. Please refer to the original papers for detailed architectural descriptions. All three models consist of four VFM stages, with output feature map resolutions of $1/4\times 1/4$, $8\times 1/8, 1/1$$16\times 1/16$, and $1/32\times 1/32$ relative to the input image resolution at stages 1 through 4, respectively. In §4, we will explore to use their first few stages as HMR and HPE encoders.
HMR2.0 [goel+2023iccv] HMR2.0-L HMR2.0-B HMR2.0-S
Encoder ViTPose-H ViTPose-L ViTPose-B ViTPose-S
(631.0 M) (303.3 M) (85.8 M) (21.7 M)
Decoder NN 6 6 3 3
$h$ 8 8 8 4
$d_{\text{hid}}$ 64 32 24 16
$d_{\text{ff}}$ 1024 512 384 128
(39.5 M) (19.1 M) (7.0 M) (2.3 M)
total (670.5M) (322.4M) (92.8M) (24.0M)
Table 1: Building on the original HMR2.0 architecture proposed by [goel+2023iccv], we introduce three scaled variants: HMR2.0-L, HMR2.0-B, and HMR2.0-S. NN denotes the number of Transformer layers, $h$ represents the number of attention heads in each cross-attention layer, $d_{\text{hid}}$ is the hidden dimension size, and $d_{\text{ff}}$ refers to the hidden dimension size of the feed-forward MLP block.
Figure 3: We explore the efficient utilization of hierarchical VFMs as encoders for HMR and HPE. (a) When using all four stages, the output feature resolution of stage 4 is $1/2\times 1/2$ relative to that of non-hierarchical VFMs. To match the expected resolution, we apply a deconvolution layer with stride 2 to upsample the features. (b) When using up to stage 3, the output resolution matches that of non-hierarchical VFMs, so we directly feed the features into the decoder without additional processing. (c) When using up to stage 2, we apply a convolution layer with stride 2 to downsample the features to the target resolution.
Up to Stage 4 (S4) Up to Stage 3 (S3) Up to Stage 2 (S2)
P F $\Phi^{\text{P,2D}}$ P $\Delta$ F $\Delta$ $\Phi^{\text{P,2D}}$ $\Delta$ P $\Delta$ F $\Delta$ $\Phi^{\text{P,2D}}$ $\Delta$
SwinPose-B 92.0 19.1 77.8 62.6 -32.0 17.3 -9.7 77.7 -0.1 5.8 -93.7 4.2 -78.2 61.2 -21.4
SwinPose-S 52.6 11.3 77.3 36.1 -31.4 10.2 -9.3 77.4 0.1 4.1 -92.2 2.8 -75.1 56.8 -26.5
SwinPose-T 31.3 6.3 76.3 14.8 -52.8 5.3 -16.6 75.9 -0.6 4.1 -86.8 2.8 -55.6 57.7 -24.4
GMFPose-B 48.3 18.3 79.7 24.9 -48.5 17.1 -6.4 79.6 -0.2 9.9 -79.6 14.1 -22.9 79.5 -0.2
GMFPose-S 24.9 6.2 77.7 19.4 -22.2 5.9 -4.6 77.4 -0.4 4.3 -82.9 2.8 -53.8 74.3 -4.4
GMFPose-T 12.8 4.6 78.3 9.7 -24.1 4.5 -3.5 78.3 0.0 3.8 -70.5 3.3 -29.2 75.3 -3.8
VMPose-B 92.8 16.4 78.5 58.6 -36.9 14.7 -10.1 78.5 0.1 6.2 -93.3 4.6 -72.0 74.7 -4.8
VMPose-S 53.2 9.7 77.8 33.9 -36.3 8.8 -9.6 77.8 0.0 4.4 -91.8 3.1 -68.6 72.3 -7.1
VMPose-T 33.3 6.0 77.7 17.0 -48.9 5.2 -13.0 77.8 0.1 4.1 -87.6 2.7 -54.7 70.6 -9.1
(a) HPE models. $\Phi^{\text{P,2D}}$ is the HPE performance score, as defined in Equation (1). A higher value of $\Phi^{\text{P,2D}}$ indicates better performance.
S4 S3 S2
P F $\Phi^{\text{M,2D}}$ P $\Delta$ F $\Delta$ $\Phi^{\text{M,2D}}$ $\Delta$ P $\Delta$ F $\Delta$ $\Phi^{\text{M,2D}}$ $\Delta$
SwinHMR2.0-B 95.5 18.0 80.2 66.1 -30.8 16.2 -10.3 79.9 -0.5 9.3 -90.2 3.1 -82.9 66.1 -17.7
SwinHMR2.0-S 52.3 10.2 80.0 35.8 -31.6 9.1 -10.2 78.9 -1.4 3.8 -92.7 1.7 -83.0 60.4 -24.5
SwinHMR2.0-T 31.0 5.2 77.7 14.5 -53.4 4.2 -19.8 75.7 -2.5 3.8 -87.7 1.7 -67.0 61.3 -21.1
GMFHMR2.0-B 52.4 17.3 82.1 29.0 -44.7 16.1 -6.8 81.5 -0.7 14.0 -73.4 13.1 -24.2 79.3 -3.4
GMFHMR2.0-S 24.8 5.1 80.6 19.3 -22.2 4.8 -5.5 79.9 -0.8 4.2 -83.2 1.8 -64.8 72.4 -10.2
GMFHMR2.0-T 13.1 3.7 80.1 10.1 -23.5 3.5 -4.3 79.0 -1.4 4.1 -68.5 2.3 -36.7 72.9 -9.0
VMHMR2.0-B 96.3 15.3 81.2 62.1 -35.5 13.6 -10.8 80.3 -1.1 9.8 -89.9 3.5 -77.2 72.9 -10.3
VMHMR2.0-S 52.9 8.6 81.0 33.6 -36.5 7.7 -10.8 80.1 -1.1 4.1 -92.3 2.0 -77.3 68.8 -15.1
VMHMR2.0-T 33.0 4.9 79.6 14.5 -56.1 4.2 -14.3 78.6 -1.4 3.8 -88.5 1.6 -66.9 67.5 -15.3
(b) HMR models for 2D pose estimation. $\Phi^{\text{M,2D}}$ indicates the HMR performance score for 2D pose estimation, as defined in Equation (2). A higher value of $\Phi^{\text{M,2D}}$ indicates better performance.
S4 S3 S2
P F $\Phi^{\text{M,3D}}$ P $\Delta$ F $\Delta$ $\Phi^{\text{M,3D}}$ $\Delta$ P $\Delta$ F $\Delta$ $\Phi^{\text{M,3D}}$ $\Delta$
SwinHMR2.0-B 95.5 18.0 55.6 66.1 -30.8 16.2 -10.3 55.5 -0.2 9.3 -90.2 3.1 -82.9 67.7 21.6
SwinHMR2.0-S 52.3 10.2 55.7 35.8 -31.6 9.1 -10.2 55.5 -0.3 3.8 -92.7 1.7 -83.0 73.8 32.7
SwinHMR2.0-T 31.0 5.2 56.8 14.5 -53.4 4.2 -19.8 57.9 2.0 3.8 -87.7 1.7 -67.0 72.6 27.8
GMFHMR2.0-B 52.4 17.3 55.0 29.0 -44.7 16.1 -6.8 55.0 0.0 14.0 -73.4 13.1 -24.2 55.4 0.8
GMFHMR2.0-S 24.8 5.1 56.0 19.3 -22.2 4.8 -5.5 56.0 0.0 4.2 -83.2 1.8 -64.8 60.7 7.9
GMFHMR2.0-T 13.1 3.7 56.1 10.1 -23.5 3.5 -4.3 56.1 0.0 4.1 -68.5 2.3 -36.7 59.7 6.5
VMHMR2.0-B 96.3 15.3 54.6 62.1 -35.5 13.6 -10.8 55.0 0.7 9.8 -89.9 3.5 -77.2 58.6 7.4
VMHMR2.0-S 52.9 8.6 55.9 33.6 -36.5 7.7 -10.8 56.1 0.3 4.1 -92.3 2.0 -77.3 62.3 11.4
VMHMR2.0-T 33.0 4.9 55.7 14.5 -56.1 4.2 -14.3 55.5 -0.4 3.8 -88.5 1.6 -66.9 64.1 15.0
(c) HMR models for 3D pose estimation. $\Phi^{\text{M,3D}}$ indicates the HMR performance score for 3D pose estimation, as defined in Equation (3). A lower value of $\Phi^{\text{M,3D}}$ indicates better performance.
Table 2: Performance comparison of HPE/HMR Models with varying encoder stage depths. In each table, P denotes the number of model parameters (in millions), and F indicates the computational cost in GFLOPs. $\Delta$ represents the relative performance change (in percentage) compared to the corresponding full-stage model (i.e., S4). Red text indicates improvement, while blue text denotes degradation compared to the full-stage model. Models highlighted with green cells are used in the following comparisons with existing methods.
3 Baseline for HMR & HPE
To ensure better self-containment, we briefly review ViTPose [xu+2022neurips] and HMR2.0 [goel+2023iccv] in this section, as they serve as baselines for HPE and HMR, respectively. Both ViTPose and HMR2.0 employ a non-hierarchical VFM, ViT [dosovitskiy+2020iclr] as their encoder. While ViTPose explores four ViT variants of different sizes, HMR2.0 utilizes only the largest variant. To establish a more comprehensive baseline, we introduce smaller ViT-based variants of HMR2.0 in § 3.2.1, using encoders inherited from the corresponding ViTPose models.
3.1 ViTPose [xu+2022neurips]
As illustrated in Figure 2 (a), ViTPose adopts a straightforward encoder-decoder architecture: a ViT serves as the encoder to generate a feature map, while the decoder comprises deconvolution layers followed by a prediction layer that outputs heatmaps corresponding to keypoints. Given an image of a person with height $H$ and width $W$ (typically, $H=256$ and $W=192$), the encoder produces features with a spatial resolution of $H/16\times W/16$. In this work, we utilize the official ViTPose repository [^1] https://github.com/ViTAE-Transformer/ViTPose, including pretrained weights, to deploy four model variants, i.e., ViTPose-H, ViTPose-L, ViTPose-B, and ViTPose-S, all of which are finetuned on the COCO dataset [lin+2014eccv]. The results, including model size, computational complexity (GFLOPs), and frames per second (FPS), are summarized in Table 3.
3.2 HMR2.0 [goel+2023iccv]
HMR2.0 also adopts an encoder-decoder architecture. As illustrated in Figure 2 (b), a ViT-based encoder produces a feature map, which is subsequently flattened and passed to a transformer-based decoder to predict the SMPL [loper+2015tog] parameters, i.e., the pose parameters $\alpha$, shape parameters $\beta$, and camera parameters $\theta$. During training, the encoder is initialized with pretrained weights from the ViTPose. Note that in [goel+2023iccv], only ViTPose-H is used as the encoder. Its architectural specifications are provided in Table 1, and the corresponding 2D and 3D pose estimation results obtained using the official code [^2] https://github.com/shubham-goel/4D-Humans are reported in Table 4.
3.2.1 HMR2.0 with Smaller ViTs
Smaller ViTPose-based encoders can be integrated into the HMR2.0 framework in a straightforward manner. Based on the encoder and decoder parameter sizes of the original HMR2.0 with the ViTPose-H encoder, we construct three additional HMR2.0 variants i.e., HMR2.0-L, HMR2.0-B and HMR2.0-S. Their architectural specifications are summarized in Table 1. Each model is initialized with pretrained weights from the corresponding ViTPose encoder and trained under the same settings as in [goel+2023iccv].
The resulting 2D and 3D pose estimation performances are summarized in Table 3 and 4. As expected, the accuracies gradually decrease as the model size becomes smaller.
4 Hierarchical VFM as HMR / HPE Encoder
In this work, we explore the potential of hierarchical VFMs for efficient HMR and HPE. To this end, our design is guided by two principles: (1) maintaining a simple architecture by avoiding complex or highly specialized modules, and (2) preserving architectural consistency with the corresponding HMR2.0 and ViTPose baselines described in §3.
As noted in §2.3, most hierarchical VFMs comprise four stages, with the output resolutions of stages 2, 3, and 4 being $2\times 2$, $1\times 1$, and $1/2\times 1/2$ relative to the output resolution of non-hierarchical VFMs, respectively. Therefore, when utilizing all four stages of a hierarchical VFM, we add a $2\times 2$ deconvolution layer to align the output resolution with that of non-hierarchical VFMs. Conversely, when using only the first two stages, we insert a convolutional layer with stride 2 after stage 2 to downsample the feature map. When using the first three stages, the output from stage 3 is directly fed into the decoder without additional processing. These configurations are illustrated in Figure 3. Notice that, to keep our modifications minimal, the number of channels in the added convolutional or deconvolutional layers is matched to the output channel size of stage 3 for each VFM.
Following this principle, we instantiate HMR and HPE models based on the HMR2.0 and ViTPose frameworks using three hierarchical VFMs: Swin [liu+2021iccv], GMF [ge+2023arxiv], and VM [liu+2024neurips]. For clarity, we adopt a consistent naming convention where each model name comprises the encoder identifier, followed by the task identifier (i.e., Pose for HPE or HMR2.0 for HMR), the encoder size (i.e., Base (B), Small (S), or Tiny (T)), and a suffix indicating the number of encoder stages used. For example, SwinPose-S-S3 denotes an HPE model that uses the small variant of Swin Transformer up to Stage 3, while VMHMR2.0-T-S2 refers to an HMR model based on the tiny variant of VMamba using up to Stage 2. For the decoder, we adopt the same architecture as ViTPose or HMR2.0, matched to the corresponding encoder size. Therefore, models such as GMFPose-B-S4, GMFPose-B-S3, and GMFPose-B-S2 share the same decoder architecture, regardless of the number of encoder stages utilized. Note that for HPE, we follow the decoder design of the ViTPose variants [xu+2022neurips]. For HMR, the decoder architecture corresponding to each encoder size (i.e., base and small variants of each VFM) is consistent with those defined in Section 3.2.1 (cf. Table 1). Additionally, for tiny encoder models, we use the same decoder architecture as that of the small variants.
Figure 4: Qualitative results of hierarchical-VFM-based HPE models.
5 Evaluation
5.1 Setting
We follow the training and evaluation protocols established by our baselines [xu+2022neurips, goel+2023iccv] for HMR and HPE. For HPE, we use the COCO dataset [lin+2014eccv] for both training and evaluation. All the hierarchical-VFM-based encoders are initialized with ImageNet-1K pre-trained weights provided by their respective official repositories [^3] https://github.com/microsoft/Swin-Transformer [^4] https://github.com/AILab-CVC/GroupMixFormer [^5] https://github.com/MzeroMiko/VMamba. HPE performance is evaluated using Average Precision (AP) and Average Recall (AR) metrics. For HMR, we train our models on a mixed dataset comprising Human3.6M [ionescu+2014tpami], MPI-INF-3DHP [mehta+20173dv], COCO [lin+2014eccv], MPII [andriluka+2014cvpr], InstaVariety [kanazawa+2019cvpr], AVA [gu+2018cvpr], and AI Challenger [wu+2019icme]. We evaluate 2D pose estimation accuracy using LSP-Extended [johnson+2011cvpr], COCO-val [lin+2014eccv], and PoseTrack-val [andriluka+2018cvpr], reporting the Percentage of Correct Keypoints (PCK) of reprojected keypoints at thresholds 0.05 and 0.1. For 3D pose accuracy, we use the 3DPW-test [kolotouros+2019iccv] and Human3.6M-val [ionescu+2014tpami] datasets, reporting both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned MPJPE (PA-MPJPE). All models are trained using 8 A100 GPUs. Inference speed, measured in frames per second (FPS), is evaluated using a single A100 GPU.
The results of all hierarchical-VFM-based HPE and HMR models (27 models in total), along with ViT-based models, are presented in Table 3 and Table 4. Based on these results, the following subsections provide a validation of our key observations and a comparison with existing lightweight models.
Figure 5: Qualitative results of hierarchical-VFM-based HMR models.
5.2 Truncated Hierarchical VFMs
Here we evaluate our key idea: leveraging only the initial few stages of hierarchical VFMs (i.e., truncated hierarchical VFMs) for efficient HPE and HMR. To enable a simplified and unified comparison, we define three aggregated performance scores: $\Phi^{\text{P,2D}}$ for 2D pose estimation by HPE, $\Phi^{\text{M,2D}}$ for 2D pose estimation by HMR, and $\Phi^{\text{M,3D}}$ for 3D pose estimation by HMR. Each metric is computed as the average of the respective evaluation scores across relevant datasets. Formally, the metrics are defined as follows:
$\Phi^{\text{P,2D}}$ $=\frac{1}{|\mathcal{D}^{\text{P,2D}}|}\sum_{\mathcal{D}^{\text{P,2D}}}\frac{1}{2}(\text{AP}+\text{AR}),$ (1)
$\Phi^{\text{M,2D}}$ $=\frac{1}{|\mathcal{D}^{\text{M,2D}}|}\sum_{\mathcal{D}^{\text{M,2D}}}\frac{1}{2}(\text{PCK@0.05}+\text{PCK@0.1}),$ (2)
$\Phi^{\text{M,3D}}$ $=\frac{1}{|\mathcal{D}^{\text{M,3D}}|}\sum_{\mathcal{D}^{\text{M,3D}}}\frac{1}{2}(\text{MPJPE}+\text{PA-MPJPE}),$ (3)
where $\mathcal{D}^{\text{P,2D}}$, $\mathcal{D}^{\text{M,2D}}$, and $\mathcal{D}^{\text{M,3D}}$ denote the datasets for evaluating pose estimation by HPE, 2D pose estimation by HMR, and 3D pose estimation by HMR, respectively [^6] Specifically, in this study $\mathcal{D}^{\text{HPE,2D}}$ consists of COCO-val [lin+2014eccv], $\mathcal{D}^{\text{HMR,2D}}$ consists of LSP-Extended [johnson+2011cvpr], COCO-val [lin+2014eccv], and PoseTrack-val [andriluka+2018cvpr], and $\mathcal{D}^{\text{HMR,3D}}$ consists of 3DPW-test [kolotouros+2019iccv] and Human3.6M-val [ionescu+2014tpami]. . Note that higher values of $\Phi^{\text{HPE}}$ and $\Phi^{\text{HMR,2D}}$ indicate better performance, while for $\Phi^{\text{HMR,3D}}$, lower values are preferable.
Tables 2 (a)–(c) present the model size, computational complexity, and the performance scores defined above for nine hierarchical-VFM-based models, each utilizing a different number of encoder stages. As expected, using fewer VFM stages leads to reduced model size and lower computational cost for both HPE and HMR models. Notably, models using up to stage 3 perform comparably to those employing all four stages. Surprisingly, in some cases such as 2D pose estimation by HPE and 3D pose estimation by HMR, models with only three stages even slightly outperform their full-stage counterparts. In the case of GMF-B, using just the first two stages still yields performance close to that of the full four-stage model.
These trends are also reflected in the qualitative results shown in Figure 4 and 5. Overall, the results confirm that the first two or three stages of hierarchical VFMs serve as efficient and effective encoders for both HPE and HMR. Based on these findings, we select the models highlighted in green in Table2 for comparison with existing methods.
5.3 Comparison to Existing Methods
The left panel of Figure 1 illustrates the 3D pose estimation accuracy of HMR models on the Human3.6M dataset [ionescu+2014tpami], including selected hierarchical-VFM-based models, ViT-based HMR2.0 variants, and existing approaches [lin+2021cvpr, cho+2022eccv, dou+2023iccv]. The right panel presents the 2D pose estimation performance of HPE models, namely, hierarchical-VFM-based models, small ViTPose variants, and existing methods [li+2022eccv, li+2022sensors, xu+2021cvpr], on the COCO dataset [lin+2014eccv]. In the left panel, models positioned closer to the bottom-left corner are both more accurate and computationally efficient, while in the right panel, those closer to the top-left corner are preferred. Results in these figures demonstrate that our hierarchical-VFM-based models generally achieve lower PA-MPJPE with smaller model sizes and reduced computational cost for HMR, and offer competitive accuracy for HPE. Despite their simplicity, the findings suggest that hierarchical-VFM-based models can serve as strong and efficient baselines for both HPE and HMR tasks.
6 Conclusion
In this work, we explore the use of hierarchical VFMs as encoders for HMR and HPE models. Experimental results show that utilizing only the initial few stages of these models yields HMR/HPE performance that is comparable to, or in some cases slightly better than, that of full-stage counterparts, while successfully reducing model size and computational cost. Furthermore, these truncated models demonstrate a favorable trade-off between accuracy and efficiency compared to existing lightweight alternatives.
For future work, we plan to extend our study by instantiating additional models based on other hierarchical VFMs. We also aim to validate our approach on broader HMR tasks, including full-body and multi-person HMR.
COCO [lin+2014eccv]
Param. GFLOPs FPS AP AR
ViTPose-H [xu+2022neurips] 637.2 125.9 67 79.1 84.1
ViTPose-L [xu+2022neurips] 308.5 61.6 126 78.3 83.5
ViTPose-B [xu+2022neurips] 90.0 18.5 393 75.8 81.1
ViTPose-S [xu+2022neurips] 24.3 5.6 936 73.8 79.2
SwinPose-B-S4 92.0 19.1 284 75.1 80.4
SwinPose-B-S3 62.6 17.3 307 75.1 80.3
SwinPose-B-S2 5.8 4.2 978 57.8 64.5
SwinPose-S-S4 52.6 11.3 406 74.5 80.1
SwinPose-S-S3 36.1 10.2 436 74.6 80.1
SwinPose-S-S2 4.1 2.8 1197 53.2 60.5
SwinPose-T-S4 31.3 6.3 665 73.7 79.0
SwinPose-T-S3 14.8 5.3 757 73.2 78.6
SwinPose-T-S2 4.1 2.8 1185 54.2 61.3
GMFPose-B-S4 48.3 18.3 171 77.2 82.2
GMFPose-B-S3 24.9 17.1 182 77.1 82.0
GMFPose-B-S2 9.9 14.1 207 77.2 81.9
GMFPose-S-S4 24.9 6.2 480 75.2 80.3
GMFPose-S-S3 19.4 5.9 511 74.8 80.0
GMFPose-S-S2 4.3 2.8 869 71.8 76.9
GMFPose-T-S4 12.8 4.6 445 75.8 80.9
GMFPose-T-S3 9.7 4.5 468 75.7 80.9
GMFPose-T-S2 3.8 3.3 619 72.8 77.9
VMPose-B-S4 92.8 16.4 469 75.9 81.0
VMPose-B-S3 58.6 14.7 532 76.0 81.1
VMPose-B-S2 6.2 4.6 1046 72.1 77.3
VMPose-S-S4 53.2 9.7 585 75.2 80.4
VMPose-S-S3 33.9 8.8 660 75.2 80.3
VMPose-S-S2 4.4 3.1 1234 69.6 74.9
VMPose-T-S4 33.3 6.0 951 75.2 80.3
VMPose-T-S3 17.0 5.2 1067 75.3 80.4
VMPose-T-S2 4.1 2.7 1518 67.9 73.4
Table 3: Pose estimation results of HPE models on the COCO-val dataset [lin+2014eccv]. Models shaded with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.
LSP-Extended [johnson+2011cvpr] COCO [lin+2014eccv] PoseTrack [andriluka+2018cvpr] 3DPW [kolotouros+2019iccv] Human3.6M [ionescu+2014tpami]
Param. GFLOPs FPS P@0.05 P@0.1 P@0.05 P@0.1 P@0.05 P@0.1 M PA-M M PA-M
HMR2.0 [goel+2023iccv] 670.5 125.6 62 53.3 82.4 86.1 96.2 89.7 97.7 81.3 54.3 49.7 32.1
HMR2.0-L 322.4 60.6 118 50.7 80.3 85.1 95.8 89.5 97.7 80.6 53.5 57.0 35.2
HMR2.0-B 92.8 17.3 366 42.8 74.0 82.1 94.6 85.9 96.2 80.0 52.7 61.3 38.5
HMR2.0-S 24.0 4.5 834 39.7 70.8 80.7 93.8 85.5 95.9 81.2 53.4 66.1 41.3
SwinHMR2.0-B-S4 95.5 18.0 265 45.4 76.5 82.4 94.7 86.2 96.3 82.2 54.6 51.5 34.3
SwinHMR2.0-B-S3 66.1 16.2 289 43.9 75.4 82.6 94.6 86.7 96.2 81.1 54.0 52.1 34.8
SwinHMR2.0-B-S2 9.3 3.1 881 24.8 53.5 67.0 87.2 73.3 90.6 94.2 59.1 69.1 48.3
SwinHMR2.0-S-S4 52.3 10.2 377 43.6 75.7 82.9 94.6 86.9 96.4 81.3 54.4 52.1 34.8
SwinHMR2.0-S-S3 35.8 9.1 408 41.0 73.7 82.0 94.4 86.3 96.0 80.9 53.3 52.6 35.2
SwinHMR2.0-S-S2 3.8 1.7 1075 18.9 44.7 60.4 82.8 67.9 87.6 103.5 63.9 73.3 54.7
SwinHMR2.0-T-S4 31.0 5.2 611 39.7 71.8 80.6 93.8 84.5 95.6 83.5 55.4 52.0 36.1
SwinHMR2.0-T-S4 14.5 4.2 695 36.9 67.8 78.5 92.9 83.1 94.9 83.3 54.4 55.3 38.5
SwinHMR2.0-T-S4 3.8 1.7 1076 20.2 46.3 61.1 83.5 68.6 87.9 100.3 62.0 73.9 54.1
GMFHMR2.0-B-S4 52.4 17.3 160 48.5 79.2 84.2 95.4 88.3 96.9 82.3 54.8 50.3 32.7
GMFHMR2.0-B-S3 29.0 16.1 169 46.9 78.1 83.9 95.2 88.1 96.9 80.4 53.0 53.0 33.8
GMFHMR2.0-B-S2 14.0 13.1 199 41.9 73.8 82.2 94.7 86.7 96.3 79.1 52.3 55.0 35.4
GMFHMR2.0-S-S4 24.8 5.1 425 46.0 77.3 82.7 94.8 86.7 96.3 82.8 54.6 52.1 34.7
GMFHMR2.0-S-S3 19.3 4.8 453 44.2 75.2 82.4 94.7 87.0 96.2 81.4 54.1 53.2 35.4
GMFHMR2.0-S-S2 4.2 1.8 772 31.8 62.0 75.1 91.3 80.6 93.6 85.6 55.0 59.4 41.8
GMFHMR2.0-T-S4 13.1 3.7 380 44.6 76.1 82.5 94.7 86.6 96.2 82.1 54.7 52.5 35.1
GMFHMR2.0-T-S3 10.1 3.5 417 41.7 73.4 81.9 94.3 86.6 96.1 80.3 53.6 54.1 36.4
GMFHMR2.0-T-S2 4.1 2.3 559 32.4 62.9 75.5 91.5 81.2 93.9 83.3 54.0 60.3 41.3
VMHMR2.0-B-S4 96.3 15.3 433 45.9 77.7 83.7 95.2 88.1 96.8 81.5 54.3 48.5 34.0
VMHMR2.0-B-S3 62.1 13.6 487 44.6 76.1 82.9 95.0 87.0 96.5 80.7 53.6 51.6 34.0
VMHMR2.0-B-S2 9.8 3.5 943 32.8 63.0 75.5 91.5 81.1 93.4 83.3 54.2 57.0 40.0
VMHMR2.0-S-S4 52.9 8.6 540 46.1 77.1 83.4 94.9 87.8 96.6 81.4 54.3 53.0 34.8
VMHMR2.0-S-S3 33.6 7.7 613 44.7 75.6 82.6 94.9 86.5 96.4 81.2 54.1 53.2 35.8
VMHMR2.0-S-S2 4.1 2.0 1146 27.3 56.1 71.0 89.2 77.0 92.0 87.1 55.9 60.8 45.2
VMHMR2.0-T-S4 33.0 4.9 751 43.5 74.6 82.5 94.5 86.6 96.3 82.0 54.1 51.8 34.8
VMHMR2.0-T-S3 14.5 4.2 856 41.8 72.6 81.1 94.2 85.6 95.9 80.3 53.1 53.8 34.9
VMHMR2.0-T-S2 3.8 1.6 1402 25.1 53.3 69.7 88.3 76.9 91.6 88.3 57.1 63.7 47.2
Table 4: 2D and 3D pose estimation results of HMR models across five datasets. In this table, “P” denotes PCK, “M” denotes MPJPE, and “PA-M” denotes PA-MPJPE. Models shaded with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.